AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 231     Entry time: Tue Feb 22 08:31:25 2022
Author: RSS, OF, ML 
Subject: Beam OFF, Target ON Run 54  

09:30

Beam optimisation. No beam since 09:25.

Target ON and open to the ring

YR09xxx ETarget ~1.9e-10 (saturated)

Target E4 inlet pressure ~8.8E-9 mbar

Dump S3 ~ 6.8E-10 mbar

Nozzle temperature 150 K

Run 54 started, monitoring code restarted.

 

FH and ML are optimising the beam.

Attachment 1: 9-ex-2022-02-22_09-29-46.png  157 kB  | Hide | Hide all
9-ex-2022-02-22_09-29-46.png
Attachment 2: 9-ey-2022-02-22_09-30-12.png  156 kB  | Hide | Hide all
9-ey-2022-02-22_09-30-12.png
Attachment 3: 9-xy-2022-02-22_09-31-36.png  208 kB  | Hide | Hide all
9-xy-2022-02-22_09-31-36.png
Attachment 4: 9-adc-2022-02-22_09-28-36.png  102 kB  | Hide | Hide all
9-adc-2022-02-22_09-28-36.png
Attachment 5: 9-temp-2022-02-22_09-26-47.png  113 kB  | Hide | Hide all
9-temp-2022-02-22_09-26-47.png
Attachment 6: 9-merger-2022-02-22_09-28-56.png  128 kB  | Hide | Hide all
9-merger-2022-02-22_09-28-56.png
Attachment 7: 9-leakage-2022-02-22_09-29-14.png  21 kB  | Hide | Hide all
9-leakage-2022-02-22_09-29-14.png
Attachment 8: 9-spectrum-2022-02-22_09-27-11.png  176 kB  | Hide | Hide all
9-spectrum-2022-02-22_09-27-11.png
Attachment 9: 9-correlation-2022-02-22_09-27-37.png  102 kB  | Hide | Hide all
9-correlation-2022-02-22_09-27-37.png
Attachment 10: 9-good-events-2022-02-22_09-28-07.png  103 kB  | Hide | Hide all
9-good-events-2022-02-22_09-28-07.png
ELOG V3.1.4-unknown